OpenStudy (help_people):

The table below shows the distance y, in miles, traveled by a toy car in x minutes: Time (x) (minutes) 20 30 40 50 Distance (y) (miles) 5 10 15 20 Part A: What is the most likely value of the correlation coefficient of the data in the table? Based on the correlation coefficient, describe the relationship between time and distance traveled by the toy car. (4 points) [Choose the value of correlation coefficient from 1, 0.8, 0.5, 0.02] Part B: What is the value of the slope of the graph of distance versus time, and what does the slope represent? (3 points) Part C: Does the dat

OpenStudy (help_people):

Part C: Does the data in the table represent correlation or causation? Explain your answer. (3 points)

OpenStudy (help_people):

Here's what I put: Part A:a. 10 = 4k, k = 2.5 20 = 8k, k = 2.5 again, so 2.5 is a valid correlation-coefficient. In terms of x, it's 1/2.5 =0.4 Part B: Part C: The data can only represent correlation and not causation. Time cannot cause events, but events can happen at certain times.

OpenStudy (help_people):

Im not sure for part. b @ganeshie8

ganeshie8 (ganeshie8):

Part C looks good Part A is wrong

ganeshie8 (ganeshie8):

Look at the table again. As the time is increasing, is the distance travelled increasing or decreasing ?

OpenStudy (help_people):

increasing

ganeshie8 (ganeshie8):

also, notice that it is steadily increasing, so the correlation coefficeint would be "1" no calculations are required, ok

ganeshie8 (ganeshie8):

rephrase above in your own words for part A mention that the distance travelled is increasing steadily as the time is increasing, and add anything extra that comes to your mind

OpenStudy (help_people):

Wht about the other part of the question for Part A?

ganeshie8 (ganeshie8):

what other part of part A ?

OpenStudy (help_people):

what is the most likely vaule of eh correleation coefficient of teh data....

ganeshie8 (ganeshie8):

also, notice that it is steadily increasing, so the correlation coefficeint would be "1" no calculations are required, ok

OpenStudy (help_people):

Ok

OpenStudy (help_people):

Use the functions m(x) = 5x + 4 and n(x) = 6x − 9 to complete the function operations listed below. Part A: Find (m + n)(x). Show your work. (3 points) Part B: Find (m ⋅ n)(x). Show your work. (3 points) Part C: Find m[n(x)]. Show your work. (4 points)

OpenStudy (help_people):

@ganeshie8 thsi si the one i didnt do

ganeshie8 (ganeshie8):

(m + n)(x) = m(x) + n(x) simply add up the expressions for m(x) and n(x)

OpenStudy (help_people):

so 5x+4+6x-9?

ganeshie8 (ganeshie8):

Yes, simplify

OpenStudy (help_people):

11x+13 thats the asnwer for part a?

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

4 - 9 = ?

OpenStudy (help_people):

11x-5

ganeshie8 (ganeshie8):

Looks good

OpenStudy (help_people):

i thought it was all added mesed up

ganeshie8 (ganeshie8):

for part b : (m ⋅ n)(x) = m(x) * n(x) multiply both the expressions for m(x) and n(x)

OpenStudy (help_people):

for partb i wil fo (5x+4) * (6x-9)?

ganeshie8 (ganeshie8):

Yep!

OpenStudy (help_people):

so i get 30x^2-21x-36

OpenStudy (help_people):

Then for the last one ur going to have to explin step by tep with that one @ganeshie8

ganeshie8 (ganeshie8):

Haha okay last one is tricky indeed

ganeshie8 (ganeshie8):

m(x) = 5x + 4 n(x) = 6x − 9 to compute, m[n(x)], you replace "x" by "6x-9" in m(x) : m[ `n(x)` ] = 5( `6x-9` ) + 4 = ?

OpenStudy (help_people):

so thats how i solve using that equation?

OpenStudy (help_people):

and i am solving for x?

OpenStudy (help_people):

@ganeshie8

ganeshie8 (ganeshie8):

we're not solving for anything just simplify the expression

OpenStudy (help_people):

im not sure hpw to do that w/ this one

ganeshie8 (ganeshie8):

m(x) = 5x + 4 n(x) = 6x − 9 to compute, m[n(x)], you replace "x" by "6x-9" in m(x) : m[ `n(x)` ] = 5( `6x-9` ) + 4 = 30x - 45 + 4 = 30x - 41

OpenStudy (help_people):

Ok thank you hold on just a sec the two others i fiished so ...

OpenStudy (help_people):

Part A: Factor x2y2 + 6xy2 + 8y2. Show your work. (4 points) Part B: Factor x2 + 8x + 16. Show your work. (3 points) Part C: Factor x2 − 16. Show your work. (3 points)

OpenStudy (help_people):

Part A: y^2 (x+4) (x+2) How I found the answer: x^2y^2+6xy^2+8y^2 y^2 (x^2+6x+8) y^2 (x+4) (x+2) Part B: (x+4) (x+4) or (x+4) ^2 How I found the answer: The third digit in the polynomial is 16. So I will find the factors of 16. So I get 1 *16=16, 2*8=16, 4*4=16. Then I will look at the second digit in the polynomial which is 8x. The second one is the one that when you add the two numbers you get the exact answer. 1*16=16, 2*8=16,4*4=16 was our factors of 16. 4+4=8 is for 8x. So, you know, on the spot that it has to be 4, and 4. Since the trinomial x^2+8x+16 is completely postitve. That will then give us the answer (x+4) (x+4) or (x+4) ^2 Part C: (x-4) (x+4) To find the answer: This problem is known as a perfect square. So, for your problem you can always tell it will be a perfect square. For example x^2 - 64. 8*8=64 So, the rules with these is you open two parenthesis with the number as positive and a negative. To conclude the answer for C would be (x-4)(x+4).

OpenStudy (help_people):

@ganeshie8

ganeshie8 (ganeshie8):

Looks good!